# Turbo-GS: Accelerating 3D Gaussian Fitting for High-Quality Radiance Fields

### v1

@@ Abstract
Figure 1: Turbo-GS accelerates 3DGS fitting significantly while preserving rendering quality. It proposes efficient densification strategy and innovative dilated rendering allow training on 4K images in minutes—significantly outperforming baseline methods. Notably, Turbo-GS converges on the 4K bicycle scene in just 13 minutes—over 3$\times$×faster than Taming 3DGS (40 minutes), 14$\times$ faster than 3DGS (187 minutes) and Scaffold-GS (185 minutes).
@@ 1 Introduction
Building the radiance field [31] of a scene from multiple posed RGB images has recently become an important problem in computer vision given numerous applications in photorealistic novel view synthesis (NVS) [1, 33, 15, 24, 29], 3D reconstruction [43, 18], mixed reality [7], and robotics [48, 6]. While radiance fields were initially represented implicitly using a neural network [31, 1], primitive-based explicit methods have become more popular [15, 45, 3]. In particular, 3D Gaussian Splatting (3DGS) [24], a method that uses 3D Gaussian primitives, has become the method of choice for representing radiance fields. 3DGS generates high-quality novel views by using a differentiable renderer based on Gaussian splat rasterization, achieving real time rendering rates for 1080p (1K) images on a GPU.
@@
Some prior works have recognized this problem and proposed solutions. For example, learning-based methods [11, 50, 52] estimate intermediate Gaussian initializations or directly estimate the final Gaussian positions in a feed-forward manner, enabling reconstruction in a few seconds. However, these methods are limited to a fixed number of input views and their generalization abilities are not well studied. Among learning-free methods that can handle an arbitrary number of views, the focus has been on building highly-optimized CUDA implementations [30], lightweight encodings to quantize Gaussian attributes [16], or replacing the optimizer with second-order methods for quicker convergence [20]. These methods speed up optimization, but more improvements are needed, especially for high-resolution 4K images.
@@
Figure 2: Effect of Densification Rate. This plot shows the effect of densification rate with Scaffold-GS [29] versus Turbo-GS (Ours) on the Bicycle scene [1]. Scaffold-GS with densification every 100 iterations (default, orange) takes time to converge. An aggressive version of Scaffold-GS with densification every 20 iterations (green) initially shows improved convergence, but plateaus afterward. Ours (blue) produces higher-quality reconstruction with densification every 20 iterations.
@@
We focus on accelerating the per-scene optimization of 3DGS without using any learning-based prior and without sacrificing quality. A typical 3DGS pipeline consists of several components and design choices that directly impact fitting time: (1) optimizer-related components include initialization, optimizer choice, optimization scheduler, and number of iterations, and (2) non-optimizer components include the number of Gaussians, the rasterizer, Gaussian densification and pruning, and post-processing. Previous work speeds up a subset of these components, for instance, the optimizer [20] or the backward process [30]. Further improvements appear feasible, but have not yet been explored.
@@
In this paper, we introduce Turbo-GS, a method for fitting 3DGS that is several times faster than current state-of-the-art methods while matching or surpassing the quality of novel view synthesis. While existing methods [30] focus on reducing the time for one optimization step, we focus on reaching a higher quality with fewer steps. This is challenging since previous methods require a high number of iterations with a low densification rate – the interval after which 3D Gaussians are split. As shown in Figure 2 (green), increasing the densification rate results in early gains, but it eventually plateaus. Moreover, the initialization of the Gaussians has a big impact on the number of steps needed.
@@
Targeting the aforementioned issues, we contribute a series of designs that quickly stabilize the process and allow higher densification rates. First, the existing densification method [24] relies solely on Gaussian positional error but ignores appearance, resulting in poor performance in texture-less areas. We introduce position-appearance guidance for more effective densification. Second, to balance densification and Gaussian fitting quality, we propose a convergence-aware budgeting process for adding new Gaussians. Finally, we make densification more reliable using an adaptive mechanism that selectively adds Gaussians only to the most frequently visited regions. In addition, we bootstrap optimization by initializing with an upsampled SfM point cloud [38], and use a batched training strategy to reduce gradient oscillations.
@@ 2.1 Novel View Synthesis
Novel-view synthesis has gained significant traction in recent years, with Neural Radiance Fields (NeRF) [31] emerging as a standout technique for generating highly photorealistic images. NeRF achieves this by leveraging volume rendering to optimize multi-layer perceptron (MLP) weights, but the original approach is computationally expensive and requires several hours or even days for training. To address this, subsequent works have integrated NeRF with explicit representations like voxel and feature grids [39, 23, 15, 41], hash grids [32], and point-based methods [42] to dramatically accelerate training times.
3D Gaussian Splatting (3DGS) [24] extends these developments by modeling scenes as a collection of 3D Gaussians that are projected as 2D splats and combined using alpha blending to form pixel colors. This technique has gained popularity for enabling high-quality, real-time rendering. Scaffold-GS [29] further enhanced 3DGS by introducing a hierarchical structure that aligns anchors with scene geometry. They introduced a multi-resolution error-based densification strategy that further enhances the robustness of the adaptive control of Gaussians, improving both rendering quality and memory efficiency. Meanwhile, numerous works have focused on various enhancements to 3DGS, including improved rendering quality [47, 21], faster rendering [9, 16], level-of-details [37], better surface reconstruction accuracy [46, 17], memory optimization [36], and the ability to handle large-scale scenes [25]. While 3DGS delivers high-quality results with extremely fast rendering, it faces challenges such as unpredictable storage requirements and variable fitting durations, which can hinder its effectiveness in downstream applications.
@@ 2.2 Accelerating Gaussian Splatting Fitting
Recognizing the limitation in fitting speed, some recent work has focused on accelerating fitting and improving efficiency by enhancing initialization methods [22, 35, 26, 14], optimizing Gaussian budget (number of Gaussians) allocation [12, 30], and reducing overall training time [16, 34, 20].
In terms of initialization, methods such as Rain-GS [22] and Gaussian Splatting as MCMC [26] enable effective model fitting from sub-optimal, or randomly initialized point clouds, expanding the robustness of 3DGS in diverse settings. CoherentGS [35] leverages monocular depth and dense flow correspondences to provide a well-defined set of initial 3D Gaussians, while studies exploring alternatives to SfM-based initialization [14] have shown that combining improved random initialization with NeRF’s structural guidance achieves, or even surpasses, the quality of COLMAP initialization on large-scale, challenging scenes.
In managing the Gaussian budget, Mini-Splatting [12] identifies spatial inefficiencies within Gaussian distributions and introduces strategies like blur splitting, depth reinitialization, and intersection-preserving sampling to address these redundancies. Taming 3DGS [30] extends this approach by adapting Gaussian distribution to specific use cases while prioritizing perceptual quality through a flexible, score-based framework. Efforts to reduce training time, including EAGLES [16] and Compact3D [34], use quantization to streamline storage and computation, while 3DGS-LM [20] replaces the Adam optimizer with the LM optimizer to accelerate convergence. Other works improve the optimization runtime by improving the implementation of the underlying differentiable rasterizer [8, 13, 44].
A line of research has also focused on sparse-view reconstruction using a data-driven, feed-forward approach, directly generating Gaussians in a single forward pass [2, 4, 49, 5, 10]. In this work, we do not rely on priors from large foundation models, as they have limited capability to accommodate densely captured views and generally cannot handle high-resolution scene modeling.
@@ 3 Prelimilaries
In this section, we present a brief overview of 3DGS and related methods. Numerous previous works have tried to improve 3DGS along many dimensions, e.g., 2DGS [21] for geometry, Mip-Splatting [47] for antialiasing and Scaffold-GS [29] for connecting the explicit and implicit 3D representation. We choose to build Turbo-GS on top of Scaffold-GS as it is already a step towards improved structure, compactness and efficiency, potentially benefiting our goal.
@@
3DGS [24] models scene geometry as a set of 3D Gaussian primitives. Each 3D Gaussian is formulated as:
@@
Scaffold-GS [29] introduces a hierarchical, structured approach to model a scene by introducing anchor points. Each visible anchor at point $x_{v}$ spawns $k$ neural Gaussians calculated by:
@@
where $\hat{f}_{v}$ is the anchor feature, $\Delta_{vc}$ is the relative viewing distance, and $\vec{d}_{vc}$ is the direction between the camera and anchor. Neural Gaussians are then rasterized following [24]. Scaffold-GS initializes anchor points from SfM and optimizes learnable parameters and the MLPs using a loss function that combines $\mathcal{L}_{1}$, SSIM loss, and volume regularization. Anchors with low opacity values are pruned while new anchors are added where the gradient of neural Gaussians exceeds a predefined threshold. In our method, we adopt Scaffold-GS because its structured management of Gaussians provides a more stable scene structure that is suitable for frequent Gaussian operations.
@@ 4 Turbo-GS: Accelerated 3D Gaussian Fitting
Our goal is to build a method for fitting 3D Gaussians that is significantly faster than existing work without compromising radiance field quality. Unlike other methods that aim to minimize the footprint of each optimization step [20, 30], our key idea is to reach high quality with fewer steps by increasing the densification rate (Figure 2). We first describe our strategy for effective densification at a high rate (Section 4.1), followed by our initialization and dilated-rendering strategy for high resolution radiance fields.
@@ 4.1 Effective & Frequent Densification
Previous works [29, 24] split Gaussians in areas with strong positional gradients. However, this strategy fails in texture-less regions due to vanishing gradients, as shown in Figure 4, indicating that position-based densification strategy is suboptimal. To find a better densification criteria, we analyze the contribution of gradients from other Gaussian properties like covariance, opacity, and color. As shown in Figure 3, position-based densification exhibits slow convergence in the early stages, primarily due to local minima in texture-less regions. Although color-based densification converges faster, it struggles to accurately capture scene structure, resulting in overfitting to training views and poor convergence in later stages. Furthermore, we find that opacity and covariance-based densification strategies contribute little, with opacity-based densification even producing floaters. Thus, we opt to use a combination of color and position gradients as the criteria for our densification strategy, which also help effectively decouples scene appearance from geometry.
@@
Based on the above analysis, we design a densification method that combines the positional and appearance errors. We threshold the position and appearance gradient with $\tau_{\text{position}}$ and $\tau_{\text{color}}$ respectively to determine whether densification is needed. We observe that the color gradient has a smaller numerical range, so we set $\tau_{\text{color}}=0.01*\tau_{\text{position}}$. Then, we follow Scaffold-GS’s [29] design to add points with multi-scale voxels. Since appearance-based densification tends to overfit, we only activate it with a probability of 20% to supplement the position-based densification. We find that the color branch improves reconstruction even if it is not frequently activated. To enhance the densification reliability, considering that the gradient from low opacity Gaussian are sensitive, we introduce an opacity-based masking strategy that excludes low-opacity Gaussians from contributing to densification decisions.
Figure 5: Loss analysis with power function fitting. For all scenes, the $log$(loss) is linear to the $log$(iterations) after the initial stage. Thus, the relation between iteration and convergence follows a power function. We design a power-law-based adaptive budget schedule based on these insights.
@@
To create such a budget, we take a closer look at how the reconstruction process converges. We calculated the log value of the training loss and iterations in Figure 5, and observed that after the initial stage, the $log(\text{loss})$ is almost perfectly linear to the $log(\text{iterations})$ in every scene (demonstrated on MipNeRF360 [1]). This means that the relation between fitting iterations and model convergence follows a power function, which can be ascribed to the use of MSE loss [28] and L1 loss. Based on this finding, we design a power-law-based adaptive budget schedule.
@@ 5.1 Implementation Details
We build Turbo-GS on top Scaffold-GS [29], and we use the Taming 3DGS’s [30] optimized CUDA kernel for backward propagation. All of our experiments are trained for 10K iterations unless otherwise stated. In the first 3K iterations, we conduct densification for every 20 steps after 300 warm-up steps. All evaluations were conducted on a single NVIDIA RTX A6000 GPU.
@@ 5.2 Dataset and Baselines
We evaluate our method on all nine scenes of MipNeRF-360 [1] and two scenes of Deep Blending [19]. For MipNeRF-360, we downsample outdoor scenes by four and indoor scenes by two, as described in 3DGS [24]. For other scenes, we use the original resolution.
Metrics. We use standard quality metrics Peak Signal-to-Noise Ratio (PSNR), Structural Similarity (SSIM) [40] and Learned Perceptual Image Patch Similarity (LPIPS) [51]. Further, we also report the optimization time, memory and peak number of Gaussian primitives (#$\#G$) during training.
Baselines. We compare our method with popular novel-view synthesis methods: 3DGS [24], ScaffoldGS [29] and MipNeRF-360 [1]. Since our goal is to speed up fitting, we also compare with works with similar goals such as Taming 3DGS [30], Mini-Splatting[12] and EAGLES [16]. We use the default settings in the paper for these baselines.

### v2

@@ Abstract
[…]
@@ 1 Introduction
Building the radiance field [Mildenhall2020NeRF] of a scene from multiple posed RGB images has recently become an important problem in computer vision given numerous applications in photorealistic novel view synthesis (NVS) [mip360, instantngp, plenoxels, 3dgs, scaffoldgs], 3D reconstruction [yariv2023bakedsdf, guedon2024sugar], mixed reality [deng2022fov], and robotics [ze2023gnfactor, dai2023graspnerf]. While radiance fields were initially represented implicitly using a neural network [Mildenhall2020NeRF, mip360], primitive-based explicit methods have become more popular [plenoxels, plenoctrees, tensorf]. In particular, 3D Gaussian Splatting (3DGS) [3dgs], a method that uses 3D Gaussian primitives, has become the method of choice for representing radiance fields. 3DGS generates high-quality novel views by using a differentiable renderer based on Gaussian splat rasterization, achieving real time rendering rates for 1080p (1K) images on a GPU.
@@
Some prior works have recognized this problem and proposed solutions. For example, learning-based methods [instantsplat, gslrm2024, longlrm] estimate intermediate Gaussian initializations or directly estimate the final Gaussian positions in a feed-forward manner, enabling reconstruction in a few seconds. However, these methods are limited to a fixed number of input views and their generalization abilities are not well studied. Among learning-free methods that can handle an arbitrary number of views, the focus has been on building highly-optimized CUDA implementations [taming3dgs], lightweight encodings to quantize Gaussian attributes [Girish2023EAGLESEA], or replacing the optimizer with second-order methods for quicker convergence [Hllein20243DGSLMFG]. These methods speed up optimization, but more improvements are needed, especially for high-resolution 4K images.
@@
Figure 2: Effect of Densification Rate. This plot shows the effect of densification rate with Scaffold-GS [scaffoldgs] versus Turbo-GS (Ours) on the Bicycle scene [mip360]. Scaffold-GS with densification every 100 iterations (default, orange) takes time to converge. An aggressive version of Scaffold-GS with densification every 20 iterations (green) initially shows improved convergence, but plateaus afterward. Ours (blue) produces higher-quality reconstruction with densification every 20 iterations.
@@
We focus on accelerating the per-scene optimization of 3DGS without using any learning-based prior and without sacrificing quality. A typical 3DGS pipeline consists of several components and design choices that directly impact fitting time: (1) optimizer-related components include initialization, optimizer choice, optimization scheduler, and number of iterations, and (2) non-optimizer components include the number of Gaussians, the rasterizer, Gaussian densification and pruning, and post-processing. Previous work speeds up a subset of these components, for instance, the optimizer [Hllein20243DGSLMFG] or the backward process [taming3dgs]. Further improvements appear feasible, but have not yet been explored.
@@
In this paper, we introduce Turbo-GS, a method for fitting 3DGS that is several times faster than current state-of-the-art methods while matching or surpassing the quality of novel view synthesis. While existing methods [taming3dgs] focus on reducing the time for one optimization step, we focus on reaching a higher quality with fewer steps. This is challenging since previous methods require a high number of iterations with a low densification rate – the interval after which 3D Gaussians are split. As shown in Figure 2 (green), increasing the densification rate results in early gains, but it eventually plateaus. Moreover, the initialization of the Gaussians has a big impact on the number of steps needed.
@@
Targeting the aforementioned issues, we contribute a series of designs that quickly stabilize the process and allow higher densification rates. First, the existing densification method [3dgs] relies solely on Gaussian positional error but ignores appearance, resulting in poor performance in texture-less areas. We introduce position-appearance guidance for more effective densification. Second, to balance densification and Gaussian fitting quality, we propose a convergence-aware budgeting process for adding new Gaussians. Finally, we make densification more reliable using an adaptive mechanism that selectively adds Gaussians only to the most frequently visited regions. In addition, we bootstrap optimization by initializing with an upsampled SfM point cloud [colmap], and use a batched training strategy to reduce gradient oscillations.
@@ 2.1 Novel View Synthesis
Novel-view synthesis has gained significant traction in recent years, with Neural Radiance Fields (NeRF) [Mildenhall2020NeRF] emerging as a standout technique for generating highly photorealistic images. NeRF achieves this by leveraging volume rendering to optimize multi-layer perceptron (MLP) weights, but the original approach is computationally expensive and requires several hours or even days for training. To address this, subsequent works have integrated NeRF with explicit representations like voxel and feature grids [Sun2021DirectVG, Karnewar2022ReLUFT, plenoxels, Xu2023GridguidedNR], hash grids [Mller2022InstantNG], and point-based methods [Xu2022PointNeRFPN] to dramatically accelerate training times.
3D Gaussian Splatting (3DGS) [3dgs] extends these developments by modeling scenes as a collection of 3D Gaussians that are projected as 2D splats and combined using alpha blending to form pixel colors. This technique has gained popularity for enabling high-quality, real-time rendering. Scaffold-GS [scaffoldgs] further enhanced 3DGS by introducing a hierarchical structure that aligns anchors with scene geometry. They introduced a multi-resolution error-based densification strategy that further enhances the robustness of the adaptive control of Gaussians, improving both rendering quality and memory efficiency. Meanwhile, numerous works have focused on various enhancements to 3DGS, including improved rendering quality [Yu2023MipSplattingA3, Huang20242DGS], faster rendering [Fan2023LightGaussianU3, Girish2023EAGLESEA], level-of-details [ren2024octree], better surface reconstruction accuracy [Yu2024GSDF3M, Guedon2023SuGaRSG], memory optimization [Papantonakis2024ReducingTM], and the ability to handle large-scale scenes [Kerbl2024AH3]. While 3DGS delivers high-quality results with extremely fast rendering, it faces challenges such as unpredictable storage requirements and variable fitting durations, which can hinder its effectiveness in downstream applications.
@@ 2.2 Accelerating Gaussian Splatting Fitting
Recognizing the limitation in fitting speed, some recent work has focused on accelerating fitting and improving efficiency by enhancing initialization methods [Jung2024RelaxingAI, Paliwal2024CoherentGSSN, Kheradmand20243DGS, Foroutan2024EvaluatingAT], optimizing Gaussian budget (number of Gaussians) allocation [fang2024mini, taming3dgs], and reducing overall training time [Girish2023EAGLESEA, Navaneet2023CompGSSA, Hllein20243DGSLMFG].
In terms of initialization, methods such as Rain-GS [Jung2024RelaxingAI] and Gaussian Splatting as MCMC [Kheradmand20243DGS] enable effective model fitting from sub-optimal, or randomly initialized point clouds, expanding the robustness of 3DGS in diverse settings. CoherentGS [Paliwal2024CoherentGSSN] leverages monocular depth and dense flow correspondences to provide a well-defined set of initial 3D Gaussians, while studies exploring alternatives to SfM-based initialization [Foroutan2024EvaluatingAT] have shown that combining improved random initialization with NeRF’s structural guidance achieves, or even surpasses, the quality of COLMAP initialization on large-scale, challenging scenes.
In managing the Gaussian budget, Mini-Splatting [fang2024mini] identifies spatial inefficiencies within Gaussian distributions and introduces strategies like blur splitting, depth reinitialization, and intersection-preserving sampling to address these redundancies. Taming 3DGS [taming3dgs] extends this approach by adapting Gaussian distribution to specific use cases while prioritizing perceptual quality through a flexible, score-based framework. Efforts to reduce training time, including EAGLES [Girish2023EAGLESEA] and Compact3D [Navaneet2023CompGSSA], use quantization to streamline storage and computation, while 3DGS-LM [Hllein20243DGSLMFG] replaces the Adam optimizer with the LM optimizer to accelerate convergence. Other works improve the optimization runtime by improving the implementation of the underlying differentiable rasterizer [Durvasula2023DISTWARFD, Feng2024FlashGSE3, Ye2024gsplatAO].
A line of research has also focused on sparse-view reconstruction using a data-driven, feed-forward approach, directly generating Gaussians in a single forward pass [Charatan2023PixelSplat3G, Chen2024LaRaEL, Zhang2024GSLRMLR, Chen2024MVSplatE3, Fan2024InstantSplatUS]. In this work, we do not rely on priors from large foundation models, as they have limited capability to accommodate densely captured views and generally cannot handle high-resolution scene modeling.
@@ 3 Prelimilaries
In this section, we present a brief overview of 3DGS and related methods. Numerous previous works have tried to improve 3DGS along many dimensions, e.g., 2DGS [Huang20242DGS] for geometry, Mip-Splatting [Yu2023MipSplattingA3] for antialiasing and Scaffold-GS [scaffoldgs] for connecting the explicit and implicit 3D representation. We choose to build Turbo-GS on top of Scaffold-GS as it is already a step towards improved structure, compactness and efficiency, potentially benefiting our goal.
@@
3DGS [3dgs] models scene geometry as a set of 3D Gaussian primitives. Each 3D Gaussian is formulated as:
@@
Scaffold-GS [scaffoldgs] introduces a hierarchical, structured approach to model a scene by introducing anchor points. Each visible anchor at point $x_{v}$ spawns $k$ neural Gaussians calculated by:
@@
where $\hat{f}_{v}$ is the anchor feature, $\Delta_{vc}$ is the relative viewing distance, and $\vec{d}_{vc}$ is the direction between the camera and anchor. Neural Gaussians are then rasterized following [3dgs]. Scaffold-GS initializes anchor points from SfM and optimizes learnable parameters and the MLPs using a loss function that combines $\mathcal{L}_{1}$, SSIM loss, and volume regularization. Anchors with low opacity values are pruned while new anchors are added where the gradient of neural Gaussians exceeds a predefined threshold. In our method, we adopt Scaffold-GS because its structured management of Gaussians provides a more stable scene structure that is suitable for frequent Gaussian operations.
@@ 4 Turbo-GS: Accelerated 3D Gaussian Fitting
Our goal is to build a method for fitting 3D Gaussians that is significantly faster than existing work without compromising radiance field quality. Unlike other methods that aim to minimize the footprint of each optimization step [Hllein20243DGSLMFG, taming3dgs], our key idea is to reach high quality with fewer steps by increasing the densification rate (Figure 2). We first describe our strategy for effective densification at a high rate (Section 4.1), followed by our initialization and dilated-rendering strategy for high resolution radiance fields.
@@ 4.1 Effective & Frequent Densification
Previous works [scaffoldgs, 3dgs] split Gaussians in areas with strong positional gradients. However, this strategy fails in texture-less regions due to vanishing gradients, as shown in Figure 4, indicating that position-based densification strategy is suboptimal. To find a better densification criteria, we analyze the contribution of gradients from other Gaussian properties like covariance, opacity, and color. As shown in Figure 3, position-based densification exhibits slow convergence in the early stages, primarily due to local minima in texture-less regions. Although color-based densification converges faster, it struggles to accurately capture scene structure, resulting in overfitting to training views and poor convergence in later stages. Furthermore, we find that opacity and covariance-based densification strategies contribute little, with opacity-based densification even producing floaters. Thus, we opt to use a combination of color and position gradients as the criteria for our densification strategy, which also help effectively decouples scene appearance from geometry.
@@
Based on the above analysis, we design a densification method that combines the positional and appearance errors. We threshold the position and appearance gradient with $\tau_{\text{position}}$ and $\tau_{\text{color}}$ respectively to determine whether densification is needed. We observe that the color gradient has a smaller numerical range, so we set $\tau_{\text{color}}=0.01*\tau_{\text{position}}$. Then, we follow Scaffold-GS’s [scaffoldgs] design to add points with multi-scale voxels. Since appearance-based densification tends to overfit, we only activate it with a probability of 20% to supplement the position-based densification. We find that the color branch improves reconstruction even if it is not frequently activated. To enhance the densification reliability, considering that the gradient from low opacity Gaussian are sensitive, we introduce an opacity-based masking strategy that excludes low-opacity Gaussians from contributing to densification decisions.
Figure 5: Loss analysis with power function fitting. For all scenes, the loglog(loss) is linear to the loglog(iterations) after the initial stage. Thus, the relation between iteration and convergence follows a power function. We design a power-law-based adaptive budget schedule based on these insights.
@@
To create such a budget, we take a closer look at how the reconstruction process converges. We calculated the log value of the training loss and iterations in Figure 5, and observed that after the initial stage, the $log(\text{loss})$ is almost perfectly linear to the $log(\text{iterations})$ in every scene (demonstrated on MipNeRF360 [mip360]). This means that the relation between fitting iterations and model convergence follows a power function, which can be ascribed to the use of MSE loss [liang2024scaling] and L1 loss. Based on this finding, we design a power-law-based adaptive budget schedule.
@@ 5.1 Implementation Details
We build Turbo-GS on top Scaffold-GS [scaffoldgs], and we use the Taming 3DGS’s [taming3dgs] optimized CUDA kernel for backward propagation. All of our experiments are trained for 10K iterations unless otherwise stated. In the first 3K iterations, we conduct densification for every 20 steps after 300 warm-up steps. All evaluations were conducted on a single NVIDIA RTX A6000 GPU.
@@ 5.2 Dataset and Baselines
We evaluate our method on all nine scenes of MipNeRF-360 [mip360] and two scenes of Deep Blending [hedman2018deep]. For MipNeRF-360, we downsample outdoor scenes by four and indoor scenes by two, as described in 3DGS [3dgs]. For other scenes, we use the original resolution.
Metrics. We use standard quality metrics Peak Signal-to-Noise Ratio (PSNR), Structural Similarity (SSIM) [wang2004image] and Learned Perceptual Image Patch Similarity (LPIPS) [zhang2018unreasonable]. Further, we also report the optimization time, memory and peak number of Gaussian primitives (#G\#G) during training.
Baselines. We compare our method with popular novel-view synthesis methods: 3DGS [3dgs], ScaffoldGS [scaffoldgs] and MipNeRF-360 [mip360]. Since our goal is to speed up fitting, we also compare with works with similar goals such as Taming 3DGS [taming3dgs], Mini-Splatting[fang2024mini] and EAGLES [Girish2023EAGLESEA]. We use the default settings in the paper for these baselines.